OpenStudy (anonymous):

Solve this differential equation dT/dt = -3/25(T-25)

OpenStudy (lalaly):

\[\frac{dT}{T-25}=-\frac{3}{25}dt\]integrate both sides\[\ln(T-25)=-\frac{3}{25}t\]\[\large{e^{\ln(T-25)}=e^{-\frac{3}{25}t}}\]

OpenStudy (wasiqss):

right lana :)

OpenStudy (lalaly):

\[\large{T-25=e^{-\frac{3}{25}t}}\]

OpenStudy (anonymous):

ah thank you ! :)

OpenStudy (lalaly):

theres a +C after e^{3/2t}

OpenStudy (lalaly):

forgot it -.- one min let me type the answer

OpenStudy (lalaly):

\[\large{T=Ce^{-\frac{3}{25}t}+25}\]

OpenStudy (anonymous):

just integrate both the sides

OpenStudy (lalaly):

because when integrating\[\ln(T-25)=-\frac{3}{25}t+C\]so when u take e of both sides\[\large{e^{\ln(T-25)}=e^{-\frac{3}{25}t+C}}\]\[\large{\(T-25)=e^{-\frac{3}{2}t}e^c}\]e^c is just a constant so u can write it as C or any letter